Trying to solve the following:
x3-12x2+44x-48=0
(thats x cubed, 12 x square, 44x, -48)
Thanks

Answer by khwang(438)

By factoring
x^3-12x^2+44x-48 = (x-4)(x-2)(x-6) = 0
The roots are 2,4 or 6.
Solution set {2,4,6}
[or use synth. division]
1-12 + 44 -48 (4
+4 -32 +48
---------------
1 -8 + 12 (2
+2 - 12
---------------
1 -6 0